OpenStudy (anonymous):

Harder differentiation: ROUND 2. Differentiate the following function ;)

OpenStudy (anonymous):

\[y=\ln(e ^{-x} + xe ^{-x})\]

OpenStudy (anonymous):

In the illustrious words of @satellite73 .....Euler

OpenStudy (anonymous):

euler?

OpenStudy (anonymous):

\[\frac{d}{dx}[\ln(f(x)]=\frac{f'(x)}{f(x)}\]

OpenStudy (anonymous):

I know how to solve it. :)

OpenStudy (anonymous):

I know how to solve it. :)

OpenStudy (anonymous):

just curious, what language is euler?

OpenStudy (anonymous):

\[\huge y'=\frac{ d }{ dx }(\ln(e ^{-x}+\frac{ x }{ e^x })\] \[\huge y' = \frac{ \frac{ d }{ dx }(e ^{-x}+\frac{ x }{ e^x } }{ e ^{-x}+e ^{-x}x }\] \[\huge y' =- \frac{ x }{ e^x \left( e ^{-x}+\frac{ x }{ e^x } \right) }\] I skipped a lot of steps because I'm too lazy to right everything out..would take too long.

OpenStudy (anonymous):

Unfortunately, I have found a different answer

OpenStudy (anonymous):

Let me guess, y' = -x/(x+1)

OpenStudy (anonymous):

more like

OpenStudy (anonymous):

\[\frac{ -xe ^{-x} }{ xe ^{-x} + e ^{-x}}\]

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

But A for effort!